OpenStudy (anonymous):

if lim x->1 [f(x)+g(x)]=6 and lim x->1 f(x)xg(x)=4 then determine lim x->1 [f(x)-g(x)]^2. Please help and thank you!

OpenStudy (anonymous):

there was a hint for this question and it says solve it as a system...

OpenStudy (anonymous):

[f(x)-g(x)]^2 expand this and you can solve it using the laws of limit.

jimthompson5910 (jim_thompson5910):

oh maybe they want you to solve the system p+q = 6 p*q = 4 where p is the function value of f(x) when x = 1 q is the function value of g(x) when x = 1

OpenStudy (anonymous):

\[\lim_{x \rightarrow 1} (f(x)+g(x)) = \lim_{x \rightarrow 1} f(x) + \lim_{x \rightarrow 1} g(x) \] This kind of limit rule can help you with the problem

OpenStudy (freckles):

I guess we are suppose to just assume lim x->1 f(x) and/or lim x->1 g(x) exists?

OpenStudy (anonymous):

That would be the only way to solve this question. You just need to break up the 3rd equation and calculate it using the 1st and the 2nd.

OpenStudy (freckles):

But we aren't given that.

OpenStudy (anonymous):

Using jim_thompson5910's example, If p+q = 6 pq = 4 we can find (p-q)^2

OpenStudy (anonymous):

how do you isolate the variables?

OpenStudy (freckles):

\[(f-g)^2=(f+g)^2-4(fg)\]

OpenStudy (freckles):

@ganeshie8 actually beat me in typing that

OpenStudy (anonymous):

how did you get that?

OpenStudy (freckles):

\[\text{ we are given } f+g \\ \text{ so we can find } (f+g)^2 \\ \text{ but } (f-g)^2 \text{ is what we are looking for } (f-g)^2=f^2+g^2-2gf \\ \text{ but } (f+g)^2=f^2+g^2+2fg \text{ so subtracting } 4gf \text{ from that will give us } \\ (f+g)^2-4fg=(f-g)^2\]

OpenStudy (freckles):

\[\lim_{x \rightarrow 1}(f(x)+g(x))=6 \\ \text{ squaring both sides } [\lim_{x \rightarrow 1} f(x)+g(x)]^2=6^2 \\ \text{ using a limit property we can write this as } \lim_{x \rightarrow 1}(f(x)+g(x))^2=6^2 \]

OpenStudy (freckles):

\[\lim_{x \rightarrow 1}(f(x)-g(x))^2=\lim_{x \rightarrow 1}([f(x)]^2+[g(x)]^2-2 f(x)g(x)) \\ \text{ but we can't use what we are given for that } f^2+g^2 \text{ part } \\ \text{ so we have \to rewrite \in terms of what we are given } \\ =\lim_{x \rightarrow 1}([f(x)]^2+[g(x)]^2+2f(x)g(x)-2f(x)g(x)-2f(x)g(x)) \\ \text{ I added \in a zero here so I could do this: } \\ =\lim_{x \rightarrow 1}((f(x)+g(x))^2-2f(x)g(x)-2f(x)g(x)) \\ \text{ combine like terms there } \\ =\lim_{x \rightarrow 1}((f(x)+g(x))^2-4f(x)g(x))\]

OpenStudy (freckles):

\[=[\lim_{x \rightarrow 1}(f(x)+g(x))]^2-4 \cdot \lim_{x \rightarrow 1}f(x)g(x)\] @esam2 look here it is easier for me to type it here and make things look prettier :p

OpenStudy (freckles):

you are given the limit inside that square thing

OpenStudy (freckles):

you are given the limit of that one product also

OpenStudy (anonymous):

if you dont mind can we start over?

OpenStudy (freckles):

\[(a-b)^2=a^2-2ab+b^2 \\ (a+b)^2=a^2+2ab+b^2 \] are you familiar with these algebraic identities?

OpenStudy (anonymous):

yes

OpenStudy (freckles):

\[\text{ \to express} (a-b)^2 \text{ \in terms of }(a+b)^2 \text{ we need \to add } 2ab \text{ but if we add } 2ab \\ \text{ we need \to also subtract } 2ab \\ \] \[(a-b)^2=a^2-2ab+b^2+2ab-2ab \\ (a-b)^2=a^2+2ab+b^2-2ab-2ab \text{ move some terms around } \\ (a-b)^2=(a+b)^2-2ab-2ab \\ (a-b)^2=(a+b)^2-4ab \text{ combined like terms }\]

OpenStudy (freckles):

If you don't like this I think I can show you one more way how to get that equation just subtract that one squared equation I gave you from the other like so: \[\text{ } \text{ } \text{ } \text{ } (a-b)^2=a^2-2ab+b^2 \\ -[(a+b)^2=a^2+2ab+b^2]\] ------------------------------ \[(a-b)^2-(a+b)^2=0-4ab+0 \\ (a-b)^2-(a+b)^2=-4ab \\ \text{ add } (a+b)^2 \text{ on both sides } (a-b)^2=-4ab+(a+b)^2 \]

OpenStudy (anonymous):

oh i see. okay i think i got that part

OpenStudy (freckles):

\[=[\lim_{x \rightarrow 1}(f(x)+g(x))]^2-4 \cdot \lim_{x \rightarrow 1}f(x)g(x) \\ =(6)^2-4(4)\]

OpenStudy (anonymous):

so the answer will be 20?

OpenStudy (freckles):

yep

OpenStudy (anonymous):

thank you so much!!!

OpenStudy (anonymous):

quick question what happen to the f(x) * g(x) =4?

OpenStudy (freckles):

we used it

OpenStudy (freckles):

\[=[\lim_{x \rightarrow 1}(f(x)+g(x))]^2-4 \cdot \lim_{x \rightarrow 1}f(x)g(x) \\ =(6)^2-\text{ } \text{ } \text{ } \text{ } \text{ } \text{ } \text{ } \text{ } \text{ } \text{ } \text{ } \text{ } \text{ } \text{ } \text{ } \text{ }4 \cdot 4 \] tried to line it up but i'm not the best at the latex code

OpenStudy (freckles):